A piano wire 0.5m long and mass 5gm is streched by a tension of 400N . What is the number of highest overtone that can e heated by a person?

Grade:12

1 Answers

Elton
7 years ago
it is a simple substitution problem.. :)  
f=?, T=400N, l=0.5m, m(=mass/length)=10-2 
as it is a stretched string the highest frequency that can be heard is the fundamental note
f=1/2l(\sqrt{T/m})
substitute the values n u get the answer....
answer(200Hz)